At a distance r1 from a point charge, the magnitude of the electric field created by the charge is 360 N/C. At a distance r2 from the charge, the field has a magnitude of 160 N/C. Find the ratio r2/r1.
Answer #1

Electric field at a point is given by:

E = k*Q/R^2

Now given that at a distance r1 from point charge

E1 = k*Q/r1^2 = 360 N/C

at a distance r2, from point charge

E2 = k*Q/r2^2 = 160 N/C

here k and Q are constant, So

dividing eq 1/eq2

360/160 = (kQ/r1^2)/(kQ/r2^2)

9/4 = (r2/r1)^2

r2/r1 = sqrt (9/4)

r2/r1 = 3/2 = 1.5

Please Upvote.